How does famous former Seattle Mariner Ichiro’s record setting year of 2004 compare with Ted
Williams year when he was the last person to hit “over .400” in 1941? Ichiro hit .372 in 2004,
while the league average was .258. The standard deviation was .04. Ted Williams hit .406 and the
league average was .267 and the standard deviation was .03. (It was very hard to get accurate
numbers - different websites give different numbers.). More specifically, this question is asking
which person had a better relative performance. All you need to know about baseball is that a
bigger batting average is better,

Answer by Boreal(15235)

Ichiro had (0.372-0.258)/0.040=0.114/0.04=2.85 sd s above the mean
Williams had (0.406-0.267)/0.030=0.139/0.030=4.63 sd s above the mean. Williams had a much better relative performance.